# Long Chain-of-Thought for Reasoning Large Language Models

### v4

@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [552, 1049, 93, 446, 990, 727, 788]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [619, 516, 1059, 642, 525, 326, 723, 459], which even much reduce the hallucinations [272, 296]. Noteworthy methods, like Self-Backtracking [909], Refiner [564], and BackMath [1015], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [835]. Further, Havrilla et al. [224] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [915] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [995], LLM2 [895], ReST-MCTS* [994] and ReARTeR [675] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [174]. He et al. [226] extend this paradigm to multi-agent scenarios, improving both reasoning and agent system performance [901, 1086]. Moreover, Yuksekgonul et al. [971] and Peng et al. [567] further expand the paradigm by enabling automatic prompt optimization driven by LLMs. This approach facilitates more generalized and automated refinement of input prompts across a range of tasks, as opposed to focusing solely on refining output results. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [1011, 874].
@@ 6.2 Internal Exploration
As noted in Chu et al. [131], Shen et al. [634], and Yang et al. [903], SFT serves as a memory process, while RL enhances generalization [340]. Specifically, SFT stabilizes the model’s output format, whereas RL improves its generalization capacity, which can increase learning efficiency by up to eight times in tasks such as mathematical reasoning [623]. Consequently, as shown in Figure 10, leading research emphasizes the role of RL and reward strategies in enhancing the exploration capabilities of LLMs without external assistance. The performance comparison is presented in Table 5.
@@ 6.2.1 RL Strategies
Recent advancements in RL strategies for exploration have led to notable improvements in various tasks, particularly in reasoning tasks [671, 350, 294, 519, 848, 979, 948, 251, 972, 602, 142].
@@
(2) Reward-based RL: Reward-model-based RL refers to approaches that use a reward model or a verifier to guide learning and decision-making in the absence of explicit rewards [1007, 164, 622, 262, 795, 815]. Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [621], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [288]. Subsequently, ReMax [415] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [676] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [631] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [248] and Liu et al. [476] further refine GRPO with REINFORCE++ and Dr. GRPO, respectively, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [725] and [1079] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [260] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [137] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards. Finally, Yi et al. [934] introduce SPPD, which employs Process Preference Learning with Dynamic Value Margin for self-training. Liang et al. [416] introduce Self-aware Weakness-driven Problem Synthesis, a reinforcement-learning method that generates challenges tailored to an RLLM’s specific weaknesses [830, 171]. By concentrating training on its most difficult aspects, the model achieves more focused and effective reasoning improvements [653]. Wang et al. [775] introduce ROLL, a method designed to support R1-level large-scale training of RLLMs, enabling the efficient exploration and optimization of reasoning paths within the Mixture-of-Experts (MOE) structure [759]. Fu et al. [176] introduce AReaL, a large-scale asynchronous reinforcement learning system for language reasoning, which enhances the efficiency and effectiveness of training RLLMs. Ma et al. [503] introduce a novel interleaved SFT and RL method that focuses on solving the hardest questions when RL can not sample any correct answer, enabling RLLMs to learn from their mistakes and improve their reasoning capabilities.
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy [279]. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [647] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [213] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [499, 160]. Furthermore, O1-CoderZhang et al. [1035], StepCoder [151], and SWE-RL [809] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation [859, 956]. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [213] introduce this concept to effectively guide reasoning and exploration [852]. Xie et al. [852] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [85, 557, 330] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [852, 926]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [136]. Kimi-K1.5 [694], Yang et al. [908] and Arora and Zanette [21] proposed Long2Short techniques, favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes [665]. Earlier in 2021, OpenAI [135] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [728, 994, 495] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding [632]. Building on this, Hou et al. [247] introduce entropy rewards and dynamic regularization to further optimize the reasoning process [111]. STeCa [741] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [694] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [592, 514, 856, 836, 988, 973, 478, 232, 870, 510, 407, 812, 1055]. Zhang et al. [1040] introduce multimodal chain-of-thought (MMCoT), while M3CoT [86] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [267, 223, 992]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [86] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, a series of work focus on optimizing the multimodal Long CoT capabilities [529, 1062]. For example, Li et al. [410] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [149] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [667] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [866] and Virgo [156] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [707, 92, 637]. Additionally, AtomThink [845] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [805] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases. This includes model decoding [466, 56, 860, 885] and RL training [845, 776, 985, 734, 276, 571, 679, 474, 414], as well as the diffusion process [504, 939, 850], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating [120, 216, 370, 122] or grounding [824] visual content holds promise for improving complex spatial reasoning tasks, particularly when logic cannot be easily conveyed through text alone [121, 666, 91, 878]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference-time scaling can be achieved through methods like RL or MCTS [822, 290], presenting an interesting avenue for future research [468].
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [188, 658, 471, 600, 168, 469, 638, 794], such as KV Cache optimization [999, 911, 464], token compression [507, 537, 960, 200, 875, 163, 651], efficient structure [293, 263, 114, 65, 235, 354, 554, 877] and dynamic reasoning patterns [758, 145, 664, 480, 367, 307, 1017, 826, 436, 449, 846, 329, 935, 719, 1023]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT [189, 1046]. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [124, 669, 24, 246, 541, 940, 467, 117]. Specifically, a series of work [694, 493, 64, 507, 537] encourage the generation of shorter reasoning processes [33, 535, 771, 186] or removing reflection signal tokens [735], minimizing redundancy and enhancing efficiency [21, 873, 476]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [218, 978, 730, 292, 375, 5, 408]. Building on these approaches, MARP [85] and DynaThink [548] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [204, 627, 1105, 145, 138, 758, 321, 465, 313, 832, 1101]. Moreover, Botta et al. [51] and Xia et al. [842] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length [1103]. More radically, Yu et al. [947] and Du et al. [153] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [222], LaTRO [73], and SoftCoT [879] transfer reasoning into continuous latent space, promoting “continuous thinking” and enabling the model to maintain multiple alternative reasoning paths [1003, 880]. Similarly, Wang et al. [780] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.

### v5

@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [578, 1091, 98, 469, 1028, 754, 818, 546]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [645, 539, 1101, 669, 549, 345, 750, 482], which even much reduce the hallucinations [289, 315]. Noteworthy methods, like Self-Backtracking [944], Refiner [590], and BackMath [1055], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [868]. Further, Havrilla et al. [238] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [950] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [1033], LLM2 [930], ReST-MCTS* [1032] and ReARTeR [703] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [186]. He et al. [240] extend this paradigm to multi-agent scenarios, improving both reasoning and agent system performance [936, 1128]. Moreover, Yuksekgonul et al. [1009] and Peng et al. [593] further expand the paradigm by enabling automatic prompt optimization driven by LLMs. This approach facilitates more generalized and automated refinement of input prompts across a range of tasks, as opposed to focusing solely on refining output results. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [1051, 908].
@@ 6.2 Internal Exploration
As noted in Chu et al. [137], Shen et al. [661], and Yang et al. [938], SFT serves as a memory process, while RL enhances generalization [359, 82]. Specifically, SFT stabilizes the model’s output format, whereas RL improves its generalization capacity, which can increase learning efficiency by up to eight times in tasks such as mathematical reasoning [650]. Consequently, as shown in Figure 10, leading research emphasizes the role of RL and reward strategies in enhancing the exploration capabilities of LLMs without external assistance. The performance comparison is presented in Table 5.
@@ 6.2.1 RL Strategies
Recent advancements in RL strategies for exploration have led to notable improvements in various tasks, particularly in reasoning tasks [699, 369, 313, 542, 882, 1017, 985, 268, 1010, 628, 150, 176, 686].
@@
(2) Reward-based RL: Reward-model-based RL refers to approaches that use a reward model or a verifier to guide learning and decision-making in the absence of explicit rewards [1046, 174, 649, 279, 825, 847, 970]. Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [648], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [306]. Subsequently, ReMax [436] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [704] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [658] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [265] and Liu et al. [499] further refine GRPO with REINFORCE++ and Dr. GRPO, respectively, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [752] and [1121] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [277] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [143] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards.
More recently, Liang et al. [439] introduce Self-aware Weakness-driven Problem Synthesis, a reinforcement-learning method that generates challenges tailored to an RLLM’s specific weaknesses [863, 183]. By concentrating training on its most difficult aspects, the model achieves more focused and effective reasoning improvements [680]. Wang et al. [805] introduce ROLL, a method designed to support R1-level large-scale training of RLLMs, enabling the efficient exploration and optimization of reasoning paths within the Mixture-of-Experts (MOE) structure [788]. Fu et al. [188] introduce AReaL, a large-scale asynchronous reinforcement learning system for language reasoning, which enhances the efficiency and effectiveness of training RLLMs. Ma et al. [526] propose a novel method combining interleaved SFT and RL to address challenging questions where RL typically fails. This approach enables RLLMs to learn from mistakes and enhance reasoning abilities. Huang et al. [297] and Fu et al. [190] further improve exploration efficiency by integrating SFT and RL with prefix sampling. Fruthermore, Yan et al. [917] and Liang et al. [437] guide RLLMs in reasoning under off-policy reinforcement learning [413, 773], improving both training sample efficiency and learning stability [559].
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy [296]. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [674] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [227] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [522, 170]. Furthermore, O1-CoderZhang et al. [1076], StepCoder [161], and SWE-RL [841] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation [893, 994]. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [227] introduce this concept to effectively guide reasoning and exploration [886]. Xie et al. [886] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. Additionally, Wu et al. [855] propose TAPO (Thought-Augmented Policy Optimization), a framework that integrates external high-level guidance (“thought patterns”) into RL, successfully balancing model exploration with external guidance. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [90, 583, 349] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [886, 962]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [142]. Kimi-K1.5 [722], Yang et al. [943] and Arora and Zanette [22] proposed Long2Short techniques, favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes [693]. Earlier in 2021, OpenAI [141] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [755, 1032, 518] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding [659, 1042]. Building on this, Hou et al. [264] introduce entropy rewards and dynamic regularization to further optimize the reasoning process [116]. STeCa [768] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [722] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [618, 537, 890, 869, 1026, 1011, 501, 246, 904, 533, 428, 844, 1097]. Zhang et al. [1081] introduce multimodal chain-of-thought (MMCoT), while M3CoT [91] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [284, 237, 1030]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [91] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, a series of work focus on optimizing the multimodal Long CoT capabilities [554, 1104, 839]. For example, Li et al. [431] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [159] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [695] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [900] and Virgo [166] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [734, 97, 664]. Additionally, AtomThink [879] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [835] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases [82]. This includes model decoding [489, 60, 894, 920] and RL training [879, 806, 1023, 761, 293, 597, 707, 497, 435], as well as the diffusion process [527, 976, 884], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating [125, 230, 390, 127] or grounding [857, 661, 149] visual content holds promise for improving complex spatial reasoning tasks [1072], particularly when logic cannot be easily conveyed through text alone [126, 694, 96, 912]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference-time scaling can be achieved through methods like RL or MCTS [854, 308], presenting an interesting avenue for future research [491, 989].
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [201, 685, 494, 626, 180, 492, 665, 824], such as KV Cache optimization [1037, 946, 487], token compression [530, 563, 998, 214, 909, 173, 678, 249, 130], efficient structure [312, 280, 119, 69, 251, 373, 580, 911, 209] and dynamic reasoning patterns [787, 154, 692, 503, 386, 326, 1057, 859, 459, 472, 880, 348, 971, 746, 1063, 153]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT [202, 1087]. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [129, 697, 25, 263, 567, 977, 490, 122]. Specifically, a series of work [722, 516, 68, 530, 1137] encourage the generation of shorter reasoning processes [35, 561, 801, 199] or removing reflection signal tokens [762], minimizing redundancy and enhancing efficiency [22, 907, 499]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [232, 1016, 757, 311, 395, 6, 429]. Building on these approaches, MARP [90] and DynaThink [574] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [218, 654, 1148, 154, 145, 787, 340, 488, 332, 865, 1144]. Moreover, Botta et al. [55] and Xia et al. [876] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length [1146]. More radically, Yu et al. [984] and Du et al. [163] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [236], LaTRO [77], and SoftCoT [913] transfer reasoning into continuous latent space, promoting “continuous thinking” and enabling the model to maintain multiple alternative reasoning paths [1041, 914]. Similarly, Wang et al. [810] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.